8 years ago, my mother was five times older than me. She turned 43 years old this year. In 6 years, how old am I?
Found 2 solutions by math_tutor2020, Alan3354:
Answer by math_tutor2020(3817)


x = your current age, in years
x-8 = your age 8 years ago
5(x-8) = your mother's age 8 years ago
5(x-8)+8 = fast forward 8 years back to the present day

The expression 5(x-8)+8 represents your mother's current age, which is 43. This leads to the equation
5(x-8)+8 = 43

Solve for x
5(x-8)+8 = 43
5x-40+8 = 43
5x-32 = 43
5x = 43+32
5x = 75
x = 75/5
x = 15
Your current age is 15 years old.
In 6 years, you will be x+6 = 15+6 = 21.

Answer: 21

Answer by Alan3354(69443)

my mother was five times older than me.
---------------------
5 times older is 6 times as old.
=====
1 time older = 2x
2 times older = 3x
etc.